\begin{thebibliography}{49}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Agrawal et~al.(2022)Agrawal, Hegselmann, Lang, Kim, and
  Sontag}]{agrawal2022}
Monica Agrawal, Stefan Hegselmann, Hunter Lang, Yoon Kim, and David Sontag.
  2022.
\newblock \href {https://doi.org/10.18653/v1/2022.emnlp-main.130} {Large
  language models are few-shot clinical information extractors}.
\newblock In \emph{Proceedings of the 2022 Conference on Empirical Methods in
  Natural Language Processing}, pages 1998--2022, Abu Dhabi, United Arab
  Emirates. Association for Computational Linguistics.

\bibitem[{Bang et~al.(2023)Bang, Cahyawijaya, Lee, Dai, Su, Wilie, Lovenia, Ji,
  Yu, Chung, Do, Xu, and Fung}]{gangda}
Yejin Bang, Samuel Cahyawijaya, Nayeon Lee, Wenliang Dai, Dan Su, Bryan Wilie,
  Holy Lovenia, Ziwei Ji, Tiezheng Yu, Willy Chung, Quyet~V. Do, Yan Xu, and
  Pascale Fung. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2302.04023} {A multitask,
  multilingual, multimodal evaluation of chatgpt on reasoning, hallucination,
  and interactivity}.
\newblock \emph{CoRR}, abs/2302.04023.

\bibitem[{Bennani-Smires et~al.(2018)Bennani-Smires, Musat, Hossmann,
  Baeriswyl, and Jaggi}]{embedrank}
Kamil Bennani-Smires, Claudiu Musat, Andreea Hossmann, Michael Baeriswyl, and
  Martin Jaggi. 2018.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/conll/conll2018.html#Bennani-SmiresM18}
  {Simple unsupervised keyphrase extraction using sentence embeddings.}
\newblock In \emph{CoNLL}, pages 221--229. Association for Computational
  Linguistics.

\bibitem[{Boudin(2018)}]{multipartiterank}
Florian Boudin. 2018.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/naacl/naacl2018-2.html#Boudin18}
  {Unsupervised keyphrase extraction with multipartite graphs.}
\newblock In \emph{NAACL-HLT (2)}, pages 667--672. Association for
  Computational Linguistics.

\bibitem[{Bougouin et~al.(2013)Bougouin, Boudin, and Daille}]{topicrank}
Adrien Bougouin, Florian Boudin, and Béatrice Daille. 2013.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/ijcnlp/ijcnlp2013.html#BougouinBD13}
  {Topicrank: Graph-based topic ranking for keyphrase extraction.}
\newblock In \emph{IJCNLP}, pages 543--551. Asian Federation of Natural
  Language Processing / ACL.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan,
  Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess,
  Clark, Berner, McCandlish, Radford, Sutskever, and Amodei}]{llm_few}
Tom~B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan,
  Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda
  Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan,
  Rewon Child, Aditya Ramesh, Daniel~M. Ziegler, Jeffrey Wu, Clemens Winter,
  Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray,
  Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford,
  Ilya Sutskever, and Dario Amodei. 2020.
\newblock Language models are few-shot learners.
\newblock NIPS'20, Red Hook, NY, USA. Curran Associates Inc.

\bibitem[{Campos et~al.(2018)Campos, Mangaravite, Pasquali, Jorge, Nunes, and
  Jatowt}]{yake}
Ricardo Campos, Vítor Mangaravite, Arian Pasquali, Alípio~Mário Jorge,
  Célia Nunes, and Adam Jatowt. 2018.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/ecir/ecir2018.html#0001MPJNJ18a} {Yake!
  collection-independent automatic keyword extractor.}
\newblock In \emph{ECIR}, volume 10772 of \emph{Lecture Notes in Computer
  Science}, pages 806--810. Springer.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova}]{bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/naacl/naacl2019-1.html#DevlinCLT19} {Bert:
  Pre-training of deep bidirectional transformers for language understanding.}
\newblock In \emph{NAACL-HLT (1)}, pages 4171--4186. Association for
  Computational Linguistics.

\bibitem[{Ding and Luo(2021)}]{attentionrank}
Haoran Ding and Xiao Luo. 2021.
\newblock Attentionrank: Unsupervised keyphrase extraction using self and cross
  attentions.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 1919--1928.

\bibitem[{Ding and Luo(2022)}]{agrank}
Haoran Ding and Xiao Luo. 2022.
\newblock \href {https://aclanthology.org/2022.aacl-main.19} {{AGR}ank:
  Augmented graph-based unsupervised keyphrase extraction}.
\newblock In \emph{Proceedings of the 2nd Conference of the Asia-Pacific
  Chapter of the Association for Computational Linguistics and the 12th
  International Joint Conference on Natural Language Processing (Volume 1: Long
  Papers)}, pages 230--239, Online only. Association for Computational
  Linguistics.

\bibitem[{Feng et~al.(2023)Feng, Tan, Zhang, Lei, and Tsvetkov}]{ld2}
Shangbin Feng, Zhaoxuan Tan, Wenqian Zhang, Zhenyu Lei, and Yulia Tsvetkov.
  2023.
\newblock {KALM}: Knowledge-aware integration of local, document, and global
  contexts for long document understanding.
\newblock In \emph{Proceedings of the 61st Annual Meeting of the Association
  for Computational Linguistics}, pages 2116--2138. Association for
  Computational Linguistics.

\bibitem[{Florescu and Caragea(2017)}]{positionrank}
Corina Florescu and Cornelia Caragea. 2017.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/acl/acl2017-1.html#FlorescuC17}
  {Positionrank: An unsupervised approach to keyphrase extraction from
  scholarly documents.}
\newblock In \emph{ACL (1)}, pages 1105--1115. Association for Computational
  Linguistics.

\bibitem[{Hasan and Ng(2014)}]{2014survey}
Kazi~Saidul Hasan and Vincent Ng. 2014.
\newblock \href {http://dblp.uni-trier.de/db/conf/acl/acl2014-1.html#HasanN14}
  {Automatic keyphrase extraction: A survey of the state of the art.}
\newblock In \emph{ACL (1)}, pages 1262--1273. The Association for Computer
  Linguistics.

\bibitem[{Hulth(2003)}]{Inspec}
Anette Hulth. 2003.
\newblock \href {http://dblp.uni-trier.de/db/conf/emnlp/emnlp2003.html#Hulth03}
  {Improved automatic keyword extraction given more linguistic knowledge.}
\newblock In \emph{EMNLP}.

\bibitem[{Jones(2004)}]{tfidf}
Karen~Spärck Jones. 2004.
\newblock \href {http://dblp.uni-trier.de/db/journals/jd/jd60.html#Jones04} {A
  statistical interpretation of term specificity and its application in
  retrieval.}
\newblock \emph{J. Documentation}, 60(5):493--502.

\bibitem[{Kim et~al.(2010)Kim, Medelyan, Kan, and Baldwin}]{SemEval}
Su~Nam Kim, Olena Medelyan, Min-Yen Kan, and Timothy Baldwin. 2010.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/semeval/semeval2010.html#KimMKB10}
  {Semeval-2010 task 5 : Automatic keyphrase extraction from scientific
  articles.}
\newblock In \emph{SemEval@ACL}, pages 21--26. The Association for Computer
  Linguistics.

\bibitem[{Kong et~al.(2023)Kong, Zhao, Chen, Li, Qin, Sun, and
  Bai}]{promptrank}
Aobo Kong, Shiwan Zhao, Hao Chen, Qicheng Li, Yong Qin, Ruiqi Sun, and Xiaoyan
  Bai. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.acl-long.545}
  {{P}rompt{R}ank: Unsupervised keyphrase extraction using prompt}.
\newblock In \emph{Proceedings of the 61st Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 9788--9801,
  Toronto, Canada. Association for Computational Linguistics.

\bibitem[{Li et~al.(2023{\natexlab{a}})Li, Feng, Radev, and Ying}]{ld1}
Irene Li, Aosong Feng, Dragomir Radev, and Rex Ying. 2023{\natexlab{a}}.
\newblock {H}i{P}ool: Modeling long documents using graph neural networks.
\newblock In \emph{Proceedings of the 61st Annual Meeting of the Association
  for Computational Linguistics}, pages 161--171. Association for Computational
  Linguistics.

\bibitem[{Li et~al.(2023{\natexlab{b}})Li, Zhuang, Hui, Qin, Lin, Jagerman,
  Wang, and Bendersky}]{li2023generate}
Minghan Li, Honglei Zhuang, Kai Hui, Zhen Qin, Jimmy Lin, Rolf Jagerman,
  Xuanhui Wang, and Michael Bendersky. 2023{\natexlab{b}}.
\newblock \href {http://arxiv.org/abs/2311.09175} {Generate, filter, and fuse:
  Query expansion via multi-step keyword generation for zero-shot neural
  rankers}.

\bibitem[{Li et~al.(2023{\natexlab{c}})Li, Zheng, and He}]{li2023unsupervised}
Zongyi Li, Xiaoqing Zheng, and Jun He. 2023{\natexlab{c}}.
\newblock \href {http://arxiv.org/abs/2009.07481} {Unsupervised summarization
  by jointly extracting sentences and keywords}.

\bibitem[{Liang et~al.(2021)Liang, Wu, Li, and Li}]{2021unsupervised}
Xinnian Liang, Shuangzhi Wu, Mu~Li, and Zhoujun Li. 2021.
\newblock \href {https://aclanthology.org/2021.emnlp-main.14} {Unsupervised
  keyphrase extraction by jointly modeling local and global context}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 155--164, Online and Punta Cana,
  Dominican Republic. Association for Computational Linguistics.

\bibitem[{Mahata et~al.(2022)Mahata, Agarwal, Gautam, Kumar, Parekh, Singla,
  Acharya, and Shah}]{longke}
Debanjan Mahata, Navneet Agarwal, Dibya Gautam, Amardeep Kumar, Swapnil Parekh,
  Yaman~Kumar Singla, Anish Acharya, and Rajiv~Ratn Shah. 2022.
\newblock {LDKP} - {A} dataset for identifying keyphrases from long scientific
  documents.
\newblock In \emph{Proceedings of the Workshop on Deep Learning for Search and
  Recommendation {(DL4SR} 2022) co-located with the 31st {ACM} International
  Conference on Information and Knowledge Management {(CIKM} 2022)}, volume
  3317 of \emph{{CEUR} Workshop Proceedings}. CEUR-WS.org.

\bibitem[{Mihalcea and Tarau(2004)}]{textrank}
Rada Mihalcea and Paul Tarau. 2004.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/emnlp/emnlp2004.html#MihalceaT04}
  {Textrank: Bringing order into text.}
\newblock In \emph{EMNLP}, pages 404--411. ACL.

\bibitem[{Nguyen and Kan(2007)}]{Nus}
Thuy~Dung Nguyen and Min-Yen Kan. 2007.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/icadl/icadl2007.html#NguyenK07} {Keyphrase
  extraction in scientific publications.}
\newblock In \emph{ICADL}, volume 4822 of \emph{Lecture Notes in Computer
  Science}, pages 317--326. Springer.

\bibitem[{Ouyang et~al.(2022)Ouyang, Wu, Jiang, Almeida, Wainwright, Mishkin,
  Zhang, Agarwal, Slama, Ray, Schulman, Hilton, Kelton, Miller, Simens, Askell,
  Welinder, Christiano, Leike, and Lowe}]{ouyang2022training}
Long Ouyang, Jeff Wu, Xu~Jiang, Diogo Almeida, Carroll~L. Wainwright, Pamela
  Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John
  Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda
  Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022.
\newblock \href {http://arxiv.org/abs/2203.02155} {Training language models to
  follow instructions with human feedback}.

\bibitem[{Pu et~al.(2023)Pu, Wang, and Demberg}]{ld3}
Dongqi Pu, Yifan Wang, and Vera Demberg. 2023.
\newblock Incorporating distributions of discourse structure for long document
  abstractive summarization.
\newblock In \emph{Proceedings of the 61st Annual Meeting of the Association
  for Computational Linguistics}, pages 5574--5590. Association for
  Computational Linguistics.

\bibitem[{Salewski et~al.(2023)Salewski, Fauth, Koepke, and
  Akata}]{salewski2023zeroshot}
Leonard Salewski, Stefan Fauth, A.~Sophia Koepke, and Zeynep Akata. 2023.
\newblock \href {http://arxiv.org/abs/2311.08396} {Zero-shot audio captioning
  with audio-language model guidance and audio context keywords}.

\bibitem[{Saxena et~al.(2020)Saxena, Mangal, and Jain}]{keygames}
Arnav Saxena, Mudit Mangal, and Goonjan Jain. 2020.
\newblock Keygames: A game theoretic approach to automatic keyphrase
  extraction.
\newblock In \emph{Proceedings of the 28th International Conference on
  Computational Linguistics}, pages 2037--2048.

\bibitem[{Song et~al.(2022{\natexlab{a}})Song, Feng, and Jing}]{hypermatch}
Mingyang Song, Yi~Feng, and Liping Jing. 2022{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2022.naacl-main.419} {Hyperbolic
  relevance matching for neural keyphrase extraction}.
\newblock In \emph{Proceedings of the 2022 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {NAACL} 2022, Seattle, WA, United States, July 10-15, 2022},
  pages 5710--5720. Association for Computational Linguistics.

\bibitem[{Song et~al.(2022{\natexlab{b}})Song, Feng, and
  Jing}]{hypersiamesenet}
Mingyang Song, Yi~Feng, and Liping Jing. 2022{\natexlab{b}}.
\newblock \href {https://doi.org/10.1145/3511808.3557538} {A preliminary
  exploration of extractive multi-document summarization in hyperbolic space}.
\newblock In \emph{Proceedings of the 31st {ACM} International Conference on
  Information {\&} Knowledge Management, Atlanta, GA, USA, October 17-21,
  2022}, pages 4505--4509. {ACM}.

\bibitem[{Song et~al.(2022{\natexlab{c}})Song, Feng, and Jing}]{icnlsp}
Mingyang Song, Yi~Feng, and Liping Jing. 2022{\natexlab{c}}.
\newblock \href {https://aclanthology.org/2022.icnlsp-1.32} {Utilizing {BERT}
  intermediate layers for unsupervised keyphrase extraction}.
\newblock In \emph{5th International Conference on Natural Language and Speech
  Processing, {ICNLSP} 2022, Trento, Italy, December 16-17, 2022}, pages
  277--281. Association for Computational Linguistics.

\bibitem[{Song et~al.(2023{\natexlab{a}})Song, Feng, and Jing}]{HISUM}
Mingyang Song, Yi~Feng, and Liping Jing. 2023{\natexlab{a}}.
\newblock \href {https://doi.org/10.1145/3543507.3583197} {Hisum: Hyperbolic
  interaction model for extractive multi-document summarization}.
\newblock In \emph{Proceedings of the {ACM} Web Conference 2023, {WWW} 2023,
  Austin, TX, USA, 30 April 2023 - 4 May 2023}, pages 1427--1436. {ACM}.

\bibitem[{Song et~al.(2023{\natexlab{b}})Song, Feng, and Jing}]{song_survey}
Mingyang Song, Yi~Feng, and Liping Jing. 2023{\natexlab{b}}.
\newblock \href {https://aclanthology.org/2023.findings-eacl.161} {A survey on
  recent advances in keyphrase extraction from pre-trained language models}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  {EACL} 2023, Dubrovnik, Croatia, May 2-6, 2023}, pages 2108--2119.
  Association for Computational Linguistics.

\bibitem[{Song et~al.(2023{\natexlab{c}})Song, Jiang, Liu, Shi, and
  Jing}]{setmatch}
Mingyang Song, Haiyun Jiang, Lemao Liu, Shuming Shi, and Liping Jing.
  2023{\natexlab{c}}.
\newblock \href {https://aclanthology.org/2023.findings-acl.156} {Unsupervised
  keyphrase extraction by learning neural keyphrase set function}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL 2023}, pages 2482--2494, Toronto, Canada. Association for Computational
  Linguistics.

\bibitem[{Song et~al.(2023{\natexlab{d}})Song, Jiang, Shi, Yao, Lu, Feng, Liu,
  and Jing}]{song2023chatgpt}
Mingyang Song, Haiyun Jiang, Shuming Shi, Songfang Yao, Shilong Lu, Yi~Feng,
  Huafeng Liu, and Liping Jing. 2023{\natexlab{d}}.
\newblock \href {http://arxiv.org/abs/2303.13001} {Is chatgpt a good keyphrase
  generator? a preliminary study}.

\bibitem[{Song et~al.(2021{\natexlab{a}})Song, Jing, Feng, Sun, and
  Xiao}]{pmlr}
Mingyang Song, Liping Jing, Yi~Feng, Zhiwei Sun, and Lin Xiao.
  2021{\natexlab{a}}.
\newblock \href {https://proceedings.mlr.press/v157/song21a.html} {Hybrid
  summarization with semantic weighting reward and latent structure detector}.
\newblock In \emph{Proceedings of The 13th Asian Conference on Machine
  Learning}, volume 157 of \emph{Proceedings of Machine Learning Research},
  pages 1739--1754. PMLR.

\bibitem[{Song et~al.(2021{\natexlab{b}})Song, Jing, and Xiao}]{kiemp}
Mingyang Song, Liping Jing, and Lin Xiao. 2021{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.215} {{I}mportance
  {E}stimation from {M}ultiple {P}erspectives for {K}eyphrase {E}xtraction}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 2726--2736, Online and Punta Cana,
  Dominican Republic. Association for Computational Linguistics.

\bibitem[{Song et~al.(2023{\natexlab{e}})Song, Liu, Feng, and Jing}]{hguke}
Mingyang Song, Huafeng Liu, Yi~Feng, and Liping Jing. 2023{\natexlab{e}}.
\newblock \href {https://aclanthology.org/2023.findings-acl.66} {Improving
  embedding-based unsupervised keyphrase extraction by incorporating structural
  information}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL 2023}, pages 1041--1048, Toronto, Canada. Association for Computational
  Linguistics.

\bibitem[{Song et~al.(2023{\natexlab{f}})Song, Liu, and Jing}]{hyperrank}
Mingyang Song, Huafeng Liu, and Liping Jing. 2023{\natexlab{f}}.
\newblock \href {https://doi.org/10.18653/v1/2023.emnlp-main.997}
  {{H}yper{R}ank: Hyperbolic ranking model for unsupervised keyphrase
  extraction}.
\newblock In \emph{Proceedings of the 2023 Conference on Empirical Methods in
  Natural Language Processing}, pages 16070--16080, Singapore. Association for
  Computational Linguistics.

\bibitem[{Song et~al.(2023{\natexlab{g}})Song, Liu, and Jing}]{diversityrank}
Mingyang Song, Huafeng Liu, and Liping Jing. 2023{\natexlab{g}}.
\newblock \href {https://doi.org/10.1145/3583780.3615141} {Improving diversity
  in unsupervised keyphrase extraction with determinantal point process}.
\newblock In \emph{Proceedings of the 32nd ACM International Conference on
  Information and Knowledge Management}, CIKM '23, page 4294–4299.
  Association for Computing Machinery.

\bibitem[{Song et~al.(2023{\natexlab{h}})Song, Xiao, and Jing}]{csl}
Mingyang Song, Lin Xiao, and Liping Jing. 2023{\natexlab{h}}.
\newblock \href {https://doi.org/10.1016/j.csl.2023.101502} {Learning to
  extract from multiple perspectives for neural keyphrase extraction}.
\newblock \emph{Comput. Speech Lang.}, 81:101502.

\bibitem[{Song et~al.(2023{\natexlab{i}})Song, Xu, Feng, Liu, and
  Jing}]{centralityrank}
Mingyang Song, Pengyu Xu, Yi~Feng, Huafeng Liu, and Liping Jing.
  2023{\natexlab{i}}.
\newblock \href {https://doi.org/10.18653/v1/2023.emnlp-main.1017} {Mitigating
  over-generation for unsupervised keyphrase extraction with heterogeneous
  centrality detection}.
\newblock In \emph{Proceedings of the 2023 Conference on Empirical Methods in
  Natural Language Processing}, pages 16349--16359, Singapore. Association for
  Computational Linguistics.

\bibitem[{Sun et~al.(2021)Sun, Liu, Xiong, Liu, and Bao}]{baseline}
Si~Sun, Zhenghao Liu, Chenyan Xiong, Zhiyuan Liu, and Jie Bao. 2021.
\newblock Capturing global informativeness in open domain keyphrase extraction.
\newblock In \emph{CCF International Conference on Natural Language Processing
  and Chinese Computing}, pages 275--287. Springer.

\bibitem[{Sun et~al.(2020)Sun, Qiu, Zheng, Wang, and Zhang}]{sifrank}
Yi~Sun, Hangping Qiu, Yu~Zheng, Zhongwei Wang, and Chaoran Zhang. 2020.
\newblock \href
  {http://dblp.uni-trier.de/db/journals/access/access8.html#SunQZWZ20}
  {Sifrank: A new baseline for unsupervised keyphrase extraction based on
  pre-trained language model.}
\newblock \emph{IEEE Access}, 8:10896--10906.

\bibitem[{Tian et~al.(2023)Tian, Li, Wang, and Gu}]{Tian_2023}
Yuanyuan Tian, Wenwen Li, Sizhe Wang, and Zhining Gu. 2023.
\newblock \href {https://doi.org/10.1111/tgis.13059} {Semantic similarity
  measure of natural language text through machine learning and a
  <scp>keyword‐aware cross‐encoder‐ranking</scp> summarizer—a case
  study using <scp>ucgis gis</scp>\&amp;t body of knowledge}.
\newblock \emph{Transactions in GIS}, 27(4):1068–1089.

\bibitem[{Wan and Xiao(2008)}]{duc2001_singlerank}
Xiaojun Wan and Jianguo Xiao. 2008.
\newblock \href {http://dblp.uni-trier.de/db/conf/aaai/aaai2008.html#WanX08}
  {Single document keyphrase extraction using neighborhood knowledge.}
\newblock In \emph{AAAI}, pages 855--860. AAAI Press.

\bibitem[{Wei et~al.(2023)Wei, Cui, Cheng, Wang, Zhang, Huang, Xie, Xu, Chen,
  Zhang, Jiang, and Han}]{Wei2023ZeroShotIE}
Xiang Wei, Xingyu Cui, Ning Cheng, Xiaobin Wang, Xin Zhang, Shen Huang, Pengjun
  Xie, Jinan Xu, Yufeng Chen, Meishan Zhang, Yong Jiang, and Wenjuan Han. 2023.
\newblock \href {https://api.semanticscholar.org/CorpusID:257050669} {Zero-shot
  information extraction via chatting with chatgpt}.
\newblock \emph{ArXiv}, abs/2302.10205.

\bibitem[{Xiong et~al.(2019)Xiong, Hu, Xiong, Campos, and Overwijk}]{xiong19}
Lee Xiong, Chuan Hu, Chenyan Xiong, Daniel Campos, and Arnold Overwijk. 2019.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/emnlp/emnlp2019-1.html#XiongHXCO19} {Open
  domain web keyphrase extraction beyond language modeling.}
\newblock In \emph{EMNLP/IJCNLP (1)}, pages 5174--5183. Association for
  Computational Linguistics.

\bibitem[{Zhang et~al.(2022)Zhang, Chen, Wang, Deng, Zhang, Li, Wang, and
  Cao}]{mderank}
Linhan Zhang, Qian Chen, Wen Wang, Chong Deng, ShiLiang Zhang, Bing Li, Wei
  Wang, and Xin Cao. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.findings-acl.34} {{MDER}ank:
  A masked document embedding rank approach for unsupervised keyphrase
  extraction}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL 2022}, pages 396--409, Dublin, Ireland. Association for Computational
  Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{ACL2023}
\usepackage{times}
\usepackage{latexsym}
\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}
\usepackage{times}
\usepackage{latexsym}
\usepackage{bm}
\usepackage{algorithm}
\usepackage{algorithmic}
\usepackage{multirow}
\usepackage{amsfonts}
\usepackage{color}
\usepackage{amsmath, mathtools}
\usepackage{microtype}

\usepackage{inconsolata}

\title{Large Language Models as Zero-Shot Keyphrase Extractors: \\A Preliminary Empirical Study}

\author{Mingyang Song, Xuelian Geng, Songfang Yao, Shilong Lu, Yi Feng, Liping Jing\thanks{\ \ Corresponding Author} \\
	Beijing Key Lab of Traffic Data Analysis and Mining \\
	Beijing Jiaotong University, Beijing, China \\
	{\tt mingyang.song@bjtu.edu.cn} \\
}

\begin{document}
\maketitle
\begin{abstract}
Zero-shot keyphrase extraction aims to build a keyphrase extractor without training by human-annotated data, which is challenging due to the limited human intervention involved. Challenging but worthwhile, zero-shot setting efficiently reduces the time and effort that data labeling takes. Recent efforts on pre-trained large language models (e.g., ChatGPT and ChatGLM) show promising performance on zero-shot settings, thus inspiring us to explore prompt-based methods. In this paper, we ask whether strong keyphrase extraction models can be constructed by directly prompting the large language model ChatGPT. Through experimental results, it is found that ChatGPT still has a lot of room for improvement in the keyphrase extraction task compared to existing state-of-the-art unsupervised and supervised models. We have released the related data\footnote{\url{https://github.com/MySong7NLPer/ChatGPT_as_Keyphrase_Extractor}}.
\end{abstract}

\section{Introduction}
Keyphrase extraction aims to extract a set of important phrases from unstructured text into structured data formats, which is a fundamental and crucial task in natural language processing \cite{2014survey,song_survey,song2023chatgpt}. Typically, keyphrase is beneficial for various applications due to its concise and precise expression \cite{li2023generate,pmlr,Tian_2023,hypersiamesenet,li2023unsupervised,HISUM,salewski2023zeroshot}. Working with an enormous amount of labeling data is always hectic, labor-intensive, and time-consuming. Hence, many organizations and companies rely on keyphrase extraction to automate manual work with zero / few-shot settings \cite{2021unsupervised, hguke, promptrank, hyperrank, agrank}.

Recent works \cite{agrawal2022,Wei2023ZeroShotIE} on large-scale pre-trained language models, such as GPT-3 \cite{llm_few}, InstructGPT \cite{ouyang2022training} and ChatGPT 2, suggest that large language models perform well in various natural language processing downstream tasks even without tuning the parameters. Therefore, it is a challenging question: Is it feasible to prompt large language models to achieve a zero-shot keyphrase extractor? Based on these clues, in this paper, we turn to ChatGPT and hypothesize that ChatGPT is born with the right abilities to deposit a zero-shot keyphrase extractor in an interactive mode.

In this paper, we mainly focus on verifying the performance of ChatGPT on four keyphrase extraction datasets and the ability of understanding long documents. Generally, long documents often necessitate complex processing strategies \cite{longke,ld1,ld2,ld3}. In the keyphrase extraction task, while it may be feasible to design improved algorithms for handling a multitude of candidate keyphrases in long documents, we believe that the effective incorporation of advanced features, especially those with background knowledge, can enhance the efficient discrimination between keyphrases and non-keyphrases, even when dealing with a vast number of candidate keyphrases. Simultaneously, the question of how to encode long documents is a topic worthy of exploration \cite{2014survey}. Consequently, we test the capacity of ChatGPT as a general large language model to handle long documents.  Through extensive experiments, it is found that ChatGPT still has a lot of room for improvement in the keyphrase extraction task compared to existing state-of-the-art unsupervised and supervised models.

\begin{table*}[t]
	\begin{center}
		\small
		\renewcommand\arraystretch{1.3}
		\renewcommand\tabcolsep{13pt}
		\begin{tabular}{l|c|c|c|c}
			\hline \hline 
			\textbf{\textsc{Test Set}} & \textbf{\textsc{Domain}}  & \textbf{\textsc{Type}} & \textbf{\textsc{\# Doc.}} & \textbf{\textsc{Avg. \# Words}} \\ \hline
			\textsc{Inspec} \cite{Inspec} &  Scientific Abstract & Short & 500 & 134.6 \\
			\textsc{DUC2001} \cite{Nus} &  News Article & Long & 308 & 847.2  \\
			\textsc{SemEval2010} \cite{SemEval} &  Scientific Article & Long & 100 & 1587.5  \\
			\hline
			\textsc{OpenKP} \cite{xiong19} &  Open Web Domain & Long & 6,616 & 900.4   \\
			\hline\hline
		\end{tabular}
	\end{center}
	\caption{\label{dataset} Statistics of used test sets. {\# Doc.} is the number of documents in the dataset.  {Avg. \# Words} is the average number of words for documents. {Present KPs (\%)} indicates the percentage of keyphrases, which are presented in the documents. {Note that this report uses all of the test data rather than sampling part from it}.}
\end{table*}
\begin{table}[t]
	\renewcommand\arraystretch{1.7}
	\begin{center}
		\scriptsize
		\begin{tabular}{p{0.5cm}p{6cm}}
			\hline\hline
			
			\multicolumn{2}{c}{\textsc{Prompts}} \\ \hline

			{\textbf{Tp}$1$}    & {\texttt{Extract keywords from this text:}} \texttt{[\textsc{Document}]} \\ \hline
			{\textbf{Tp}$2$}    & \texttt{Extract keyphrases from this text:} \texttt{[\textsc{Document}]} \\ 
			\hline\hline
			
		\end{tabular}
		\caption{\label{case} Two prompts are designed for chatting with ChatGPT to extract keyphrases from the text document.}
	\end{center}
\end{table}

\section{ChatGPT for Keyphrase Extraction}

\subsection{Evaluation Setting}
We briefly introduce the evaluation setting, which mainly includes the compared baselines, datasets, and evaluation metrics.
Note that each time a new query is made to ChatGPT, we clear conversations to avoid the influence of previous samples, which is similar to \citet{gangda}.

In this paper, we compare ChatGPT with several state-of-the-art unsupervised keyphrase extraction models: {TF-IDF} \cite{tfidf}, {YAKE} \cite{yake}, {TextRank} \cite{textrank}, {TopicRank} \cite{topicrank}, \textsc{PositionRank} \cite{positionrank}, {MultipartiteRank} \cite{multipartiterank}, {EmbedRank} \cite{embedrank}, {KeyGames} \cite{keygames}, {SIFRank} \cite{sifrank}, {AttentionRank} \cite{attentionrank}, {JointGL} \cite{2021unsupervised}, {MDERank} \cite{mderank}, {SetMatch} \cite{setmatch}, {HGUKE} \cite{hguke}, {AGRank} \cite{agrank}, {HyperRank} \cite{hyperrank}, {CentralityRank} \cite{centralityrank}, and {PromptRank} \cite{promptrank}. In addition to unsupervised models, we compare ChatGPT with two state-of-the-art supervised models: HyperMatch \cite{hypermatch} and KIEMP \cite{kiemp}. Furthermore, we also select the large language models ChatGLM-6B\footnote{\url{https://github.com/THUDM/ChatGLM-6B}} and ChatGLM2-6B\footnote{\url{https://github.com/thudm/chatglm2-6b}} for comparison. By default, the results in this report come from the ChatGPT version on 2023.03.01.

We evaluate ChatGPT and all the baselines on four datasets: \textsc{Inspec} \cite{Inspec}, \textsc{DUC2001} \cite{duc2001_singlerank}, \textsc{SemEval2010} \cite{SemEval}, and \textsc{OpenKP} \cite{xiong19}. Table~\ref{dataset} summarizes the statistics of used test sets.

Following previous studies \cite{kiemp, hypermatch, csl}, we adopt macro averaged F1@5 and F1@M to evaluate the quality of both present and absent keyphrases. When using F1@5, blank keyphrases are added to make the keyphrase number reach five if the prediction number is less than five. Similar to the previous studies \cite{icnlsp, hguke, setmatch, diversityrank}, we employ the Porter Stemmer to remove the identical stemmed keyphrases.

\subsection{Keyphrase Extraction Prompts}

ChatGPT requires human-designed prompts or instructions as guiding information to trigger its ability to achieve downstream tasks, i.e., extracting keyphrases. It should be noted that the style and quality of prompts can greatly affect the quality of ChatGPT keyphrase extraction. High-quality prompts often achieve better task performance than low-quality prompts. Figure~\ref{case} presents the prompts designed for extracting keyphrases, which is similar to \cite{song2023chatgpt}.

\subsection{Overall Performance}

As illustrated in Table~\ref{extraction}, it can be seen that ChatGPT achieves similar performance to the unsupervised method TF-IDF with simple prompts, and far surpasses the unsupervised method TextRank. Although there is a certain gap in the performance of ChatGPT compared to supervised methods HyperMatch and KIEMP, fine-tuning large models like ChatGPT through human-annotated data may make the performance far superior to existing supervised keyphrase extraction models. In addition, it is interesting that ChatGLM2 with only 6B parameters achieved comparable performance to ChatGPT.

\begin{table*}[h]
	\begin{center}
		\scriptsize
		\renewcommand\arraystretch{1.5}
		\renewcommand\tabcolsep{7pt}
		\begin{tabular}{l|ccc|ccc|ccc}
			\hline \hline 
			\multirow{2}{*}{\textbf{\textsc{Model}}} 
			& \multicolumn{3}{c|}{\textsc{\textbf{DUC2001}}} & \multicolumn{3}{c|}{\textsc{\textbf{Inspec}}} & \multicolumn{3}{c}{\textsc{\textbf{SemEval2010}}} \\
			& {F1@5} &  {F1@10} &  {F1@15} & {F1@5} &  {F1@10} &  {F1@15} & {F1@5} &  {F1@10} &  {F1@15} \\ \hline
			\multicolumn{10}{l}{\textsc{\textbf{Statistical Models}}}\\\hline
			TF-IDF \cite{tfidf} & 9.21 & 10.63 & 11.06 & 11.28 & 13.88 & 13.83 & 2.81 & 3.48 & 3.91 \\
			YAKE \cite{yake} & 12.27 & 14.37 & 14.76 & 18.08 & 19.62 & 20.11 & 11.76 & 14.4 & 15.19 \\
			\hline
			\multicolumn{10}{l}{\textsc{\textbf{Graph-based Models}}}\\\hline
			
			TextRank \cite{textrank} & 11.80 & 18.28 & 20.22 & 27.04 & 25.08 & 36.65 & 3.80 & 5.38 & 7.65 \\
			SingleRank \cite{duc2001_singlerank} & 20.43 & 25.59 & 25.70 & 27.79 & 34.46 & 36.05 & 5.90 & 9.02 & 10.58 \\
			TopicRank \cite{topicrank} & 21.56 & 23.12 & 20.87 & 25.38 & 28.46 & 29.49 & 12.12 & 12.90 & 13.54 \\
			PositionRank \cite{positionrank} & 23.35 & 28.57 & 28.60 & 28.12 & 32.87 & 33.32 & 9.84 & 13.34 & 14.33 \\
			MultipartiteRank \cite{multipartiterank} & 23.20 & 25.00 & 25.24 & 25.96 & 29.57 & 30.85 & 12.13 & 13.79 & 14.92 \\\hline
			
			\multicolumn{10}{l}{\textsc{\textbf{Embedding-based Models}}}\\\hline
			EmbedRankd2v \cite{embedrank} & 24.02 & 28.12 & 28.82 & 31.51 & 37.94 & 37.96 & 3.02 & 5.08 & 7.23 \\
			EmbedRanks2v \cite{embedrank} & 27.16 & 31.85 & 31.52 & 29.88 & 37.09 & 38.40 & 5.40 & 8.91 & 10.06 \\
			KeyGames \cite{keygames} & 24.42 & 28.28 & 29.77 & 32.12 & {40.48} & 40.94 & 11.93 & 14.35 & 14.62 \\
			SIFRank \cite{sifrank} & 24.27 & 27.43 & 27.86 & 29.11 & 38.80 & 39.59 & - & -& -\\
			SIFRank+ \cite{sifrank} & {30.88} & 33.37 & 32.24 & 28.49 & 36.77 & 38.82 & - & - & -\\
			AttentionRank \cite{attentionrank} & - & - & - & 24.45 & 32.15 & 34.49 & 11.39 & 15.12 & 16.66 \\
			JointGL \cite{2021unsupervised} & 28.62 & {35.52} & {36.29} & {32.61} & 40.17 & {41.09}& 13.02 & {19.35} & {21.72} \\
			MDERank \cite{mderank} & 23.31 & 26.65 & 26.42 & 27.85 & 34.36 & 36.40 & {13.05} & 18.27 & 20.35 \\
			SetMatch \cite{setmatch} & {31.19} & {36.34} & {38.72} & {33.54} & {40.63} & {42.11} & {14.44} & {20.79} & {24.18} \\
			HGUKE \cite{hguke} & {31.31} & {37.24} & {38.31} & {34.18}& \textbf{41.05} & \textbf{42.16} & {14.07} & {20.52}& {23.10}\\
			AGRank \cite{agrank} & - & - & - & \textbf{34.59} & 40.70 & 41.15 & 15.37 & 21.22 & 23.72 \\
			HyperRank \cite{hyperrank} & \textbf{32.68} & \textbf{39.18} & \textbf{40.21} & {33.35}& {40.79} & {42.12} & {14.79} & {21.33} & \textbf{24.20} \\
			CentralityRank \cite{centralityrank} & {31.63} & {37.77} & {38.77} & {32.99} & {40.93} & {41.73} & {15.51} & \textbf{21.39} & {23.83} \\
			PromptRank \cite{promptrank} & 27.39 & 31.59 & 31.01 & 31.73 & 37.88 & 38.17 & \textbf{17.24} & 20.66 & 21.35 \\
			\hline
			\multicolumn{10}{l}{\textsc{\textbf{LLM-based Models}}}\\\hline
			\multicolumn{1}{l|}{{ChatGPT (gpt-3.5-turbo) - \textbf{Tp}$1$}}
			& 19.29 & 23.32 & 22.98 & 22.95 & 30.57 & 33.87 & 13.25 & 15.94 & 17.12    \\
			\multicolumn{1}{l|}{{ChatGPT (gpt-3.5-turbo) - \textbf{Tp}$2$}}
			& 21.50 & 25.03 & 24.21 & 28.07 & 34.85 & 36.69 & 13.66 & 16.11 & 16.18    \\
			\multicolumn{1}{l|}{{ChatGLM-6B - \textbf{Tp}$2$}}
			& 11.57 & 11.72 & 11.09 & 14.52 & 18.85 & 20.17 & 6.69 & 8.71 & 9.14 \\
			\multicolumn{1}{l|}{{ChatGLM2-6B - \textbf{Tp}$2$}}
			& 15.08 & 15.44 & 13.96 & 25.07 & 30.09 & 31.56 & 13.12 & 13.79 & 13.94 \\
			\hline\hline
		\end{tabular}
	\end{center}
	\caption{\label{total} Performance on DUC2001, Inspec and SemEval2010 test sets. The best results are in bold.}
\end{table*}

\begin{table}[h]
	\scriptsize
	\centering
	\renewcommand\tabcolsep{7pt}
	\renewcommand\arraystretch{1.5}
	\begin{tabular}{c|ccc}
		\hline\hline
		\multirow{2}{*}{\textbf{\textsc{Model}}} & \multicolumn{3}{c}{\textsc{\textbf{OpenKP}}} \\ 
		& {F1@1}  & {F1@3} & F1@5 \\ \hline
		\multicolumn{4}{l}{\textsc{\textbf{Unsupervised Models}}} \\\hline
		\multicolumn{1}{l|}{{TF-IDF} \cite{tfidf}}
		& 19.6 & 22.3 & 19.6  \\ 
		\multicolumn{1}{l|}{{TextRank} \cite{textrank}}
		& 5.4 & 7.6 & 7.9  \\ \hline
		
		\multicolumn{4}{l}{\textsc{\textbf{PLM-based Models}}} \\\hline
		
		\multicolumn{1}{l|}{{HyperMatch} \cite{hypermatch}}
		& 36.4 & \textbf{39.4} & 33.8  \\ 
		\multicolumn{1}{l|}{{KIEMP} \cite{kiemp}}
		& \textbf{36.9} & 39.2 & \textbf{34.0}   \\
		\hline
		
		\multicolumn{4}{l}{\textsc{\textbf{LLM-based Models}}}\\\hline
		
		\multicolumn{1}{l|}{{ChatGPT (gpt-3.5-turbo) - \textbf{Tp}$1$}}
		& 16.5 & 21.1 & 17.4  \\
		\multicolumn{1}{l|}{{ChatGPT (gpt-3.5-turbo) - \textbf{Tp}$2$}}
		& 18.0 & 21.6 & 18.7  \\
		
		\multicolumn{1}{l|}{{ChatGLM-6B - \textbf{Tp}$2$}}
		& 11.5 & 8.5 & 7.1  \\
		\multicolumn{1}{l|}{{ChatGLM2-6B - \textbf{Tp}$2$}}
		& 16.0 & 11.0 & 8.6  \\
		\hline\hline
	\end{tabular}
	
	\caption{Results of keyphrase extraction on the OpenKP dataset. F1@3 is the main evaluation metric for this dataset \cite{xiong19}. The results of the baselines are reported in their corresponding papers. The best results are highlighted in bold.}
	\label{extraction}
\end{table}
\subsection{Long Document Understanding}

Generally, handling long documents poses a significant challenge in numerous natural language processing tasks, primarily due to the presence of intricate contexts within extended texts. This challenge is particularly pressing in keyphrase generation, where effectively comprehending long documents is paramount. Many of the existing keyphrase extraction baseline models, such as \cite{baseline}, \cite{kiemp}, and \cite{hypermatch}, necessitate the truncation of input documents to conform to the input constraints imposed by the underlying backbone, such as BERT \cite{bert}, resulting in a substantial loss of information. Simultaneously, extracting keyphrases that are contextually meaningful to the document is crucial in achieving semantic document understanding.

Consequently, we conduct experiments to measure the capability of ChatGPT to understand lengthy documents, as shown in Table~\ref{extraction} and Table~\ref{total}. From the results, it can be seen that compared with baselines, the extraction results of ChatGPT are not very ideal. Of course, various factors affect the performance of ChatGPT, and the results presented in this paper are only preliminary explorations. 
In addition, the extraction results of ChatGPT on the long document dataset SemEval2010 are significantly better than those on the Inspec and DUC2001 datasets, which verifies its excellent ability to understand long documents. As a next step, we plan to select longer documents, for example, documents containing approximately 4096 words, to serve as test cases.
\section{Conclusion}
In this paper, we conduct preliminary experiments and analysis on the keyphrase extraction task. Although ChatGPT has achieved good results in some natural language processing downstream tasks, there is still room for improvement in the keyphrase extraction task. Of course, due to limited available resources, this paper only conducted experimental verification under some basic settings, which may to some extent limit the performance of ChatGPT. In addition, there are many methods to optimize the extraction results of ChatGPT, such as designing more complex and high-quality prompts, constructing contextual samples as auxiliary information, and introducing supervised fine-tuning methods. 

In future research, utilizing large language models as external knowledge-assisted methods may be a hot topic in the keyphrase extraction task.

\bibliography{anthology}
\bibliographystyle{acl_natbib}
\end{document}